OpenStudy (anonymous):

Write an equation and solve each problem. 1)Twice a number plus four equals 6. What is the number? 2)Sixteen is seven plus three times a number. Find the number. 3) Find two consecutive integers whose sum is 35. 4) Find three consecutive integers whose sum is 36.

OpenStudy (anonymous):

1) 2x+4=6 2x=2 x=1. 2) 16=7+3x 3x=9 x=3

OpenStudy (anonymous):

3)x+x+1=35 2x+1=35 2x=34 x=17,x+1=18

OpenStudy (anonymous):

4)x+x+1+x+2=36 3x+3=36 3x=33 x=11, x+1=12,x+2=13